OpenStudy (anonymous):

A blue ball is thrown upward with an initial speed of 23.9 m/s, from a height of 0.7 meters above the ground. 2.9 seconds after the blue ball is thrown, a red ball is thrown down with an initial speed of 9.6 m/s from a height of 31.1 meters above the ground. The force of gravity due to the earth results in the balls each having a constant downward acceleration of 9.81 m/s2. how long after the blue ball is thrown are the two balls in the air at the same height? I have checkedsig-figs, redone the equation, used wolframalpha, and yet the question is wrong. Thanks for the help.

OpenStudy (jt950):

I get the time t=0.29734 but the time -2.9second for red ball to be thrown will be negative so i think the question might be wrong

OpenStudy (anonymous):

That is what I thought, but I did some research and even asked my teacher and he said that there was an answer to this question. I tried using the quadratic equation and setting the respective equations equal to each other, but they come out wrong.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

\[0.7+23.9(t+2.9)-\frac{ 1 }{ 2}g( t+2.9)^2=31.1-9.6t-\frac{ 1 }{ 2 }g t^2\]Solving this equation I get: t=0.46 sec, then the time since the blue ball is launched is:t=0.46+2.9=3.36 sec

OpenStudy (jt950):

Sorry I was mistaken that red ball to be thrown "up" LOL There is solution for this answer